OpenStudy (anonymous):

Solve for x in 4 ^2x-3 = 16?

OpenStudy (anonymous):

\[4^{2x-3} =16\]

OpenStudy (mimi_x3):

Ok lets convert 16 into a base of 4 we know that \( 4^2=16\) So \[4^{2x-3}=4^2\]

OpenStudy (mimi_x3):

So if \[4^{2x-3}=4^2\] Then obviously the 2 exponents equal each other so 2x-3=2 Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

SO now solve for x

OpenStudy (anonymous):

x = -1/2?

OpenStudy (mimi_x3):

2x-3=2 2x=5 ....

OpenStudy (anonymous):

oh sorry i messed up on the charge and then you divide by two so x would be 5/2

OpenStudy (mimi_x3):

yup yup Medal +Fan

OpenStudy (anonymous):

okay can you help me with another one?

OpenStudy (mimi_x3):

Sure go for it

OpenStudy (anonymous):

\[2^{x^2+2x-3} = 1\]

OpenStudy (mimi_x3):

Ok we know that \(y^0=1\) Meaning any number to the power of 0 equals 1 The exception to this rule is \(0^0=0\) So now in order to solve for x we want both sides to have the same base so how can we write 1 with a base of 2????? Well \(2^0=1\) So \[2^{x^2+2x-3}=2^0\] Can you take it from here?

OpenStudy (anonymous):

so would it be x^2 + 2x -3

OpenStudy (mimi_x3):

\[x^2+2x-3=????\]

OpenStudy (anonymous):

2?

OpenStudy (mimi_x3):

hmmm we are equating the exponents of either side of the equal sign since they both have the same base - a base of 2 So noooo it doesnt equal 2

OpenStudy (anonymous):

so then its 0?

OpenStudy (mimi_x3):

Yes \[x^2+2x-3=0\] We now can factor and solve for x We will have 2 solutions for x

OpenStudy (anonymous):

do we factor it?

OpenStudy (anonymous):

would it be (x-1)(x-2)

OpenStudy (mimi_x3):

We are dealing with quadratics sooo ... Im lazy to start explaining quadratics rn You familiar with this ??? (x-s)(x-r)=0 So x=s and r

OpenStudy (mimi_x3):

Pretty close

OpenStudy (mimi_x3):

ax^2+bx+c=0 a=1 b=2 c=-3 We need two numbers that when you add them equal 2 and when you multiply them you get -3

OpenStudy (abb0t):

Why don't you take the natural log of both sides to make it easier.

OpenStudy (mimi_x3):

ya it wld buttt ... not sure what this b!tch of a hs teacher wants

OpenStudy (mimi_x3):

Im almost positive they want same bases cuzz it keeps working out perfectly

OpenStudy (anonymous):

i think they actually want the natural log but i don't care i just want the answer i still have 20 more questions to do

OpenStudy (anonymous):

lol

OpenStudy (abb0t):

Knew it. You're a cheater.

OpenStudy (mimi_x3):

Ohhh so take natural log if ya want .... But ya u shld know how to factor lmao

OpenStudy (abb0t):

This is a place to help, not to give answers!

OpenStudy (mimi_x3):

lmaooo abbot stopppppp I need to get to beddddddddd

OpenStudy (anonymous):

i just get confused on which you add to and which you multiply too.

OpenStudy (anonymous):

how am i a cheater.. I'm doing the work out right? the that doesn't make me a cheater a cheater would be googling the fluttering answers, you don't like me asking for help? then block me and get off my page

OpenStudy (anonymous):

so it would be (x+1)(x-3)

OpenStudy (mimi_x3):

Ok so any 2 numbers that when you add equal 2 and when you multiply equal -3 SO what are the factors of -3 ???? Only 1 and 3 obv so -1*3=-3 and 3-1=2

OpenStudy (anonymous):

so just switch the positive and negative around from 1 to 3 right?

OpenStudy (mimi_x3):

umm no its (x-1)(x+3)=0 Hence x=1 or -3

OpenStudy (anonymous):

yeah that what I meant because before i put (x+1)(x-3)

OpenStudy (mimi_x3):

ohh gotcha so yaaaa u are correct

OpenStudy (mimi_x3):

FAN ME PLSSSS

OpenStudy (anonymous):

thank you for your help and I will

OpenStudy (anonymous):

I fan'd you (:

OpenStudy (anonymous):

goodnight!

OpenStudy (mimi_x3):

Thanks good night :D